Print

Print


 Ticket Number: HEA-NOC/19990618-01         Ticket Status: OPEN/CLOSE
 Ticket Type:   Scheduled                   Ticket Source: HEA-NOC
 Ticket Scope:  Site                        Site/Line:     Avoca.heanet.ie
 Ticket Owner:  [log in to unmask]             Problem Fixer: HEA-NOC
 Ticket Opened: 19990618  12:56:05 UTC      Problem Start: 19990618  02:00 UTC+1
 Ticket Closed: 19990618  12:56:05 UTC      Problem End:   19990618  13:50 UTC+1


Problem Description:

Cron processes (sendmail + MRTG) on avoca.heanet.ie went wild causing
load avg to go through the roof.

Affected:

www, listserv, ftp.heanet.ie.

Actions:

Killed errant processes this morning and rebooted the machine.
MRTG had to be restored from backup and is not yet monitoring all
links. LISTSERV is operating and sendmail is clearing the backlog.

Time to Fix:

Approx 11 hours.

 Additional Notes:

Avoca.heanet.ie will shortly undergo an upgrade to Digital UNIX 4.0D
which will allow us to solve most of its current problems. A backup
server for LISTSERV will be in operation and WWW/Database services
will be migrated to another machine. The upgrade will be scheduled
separately.